Grade 12th passMechanics

a particle moves rectilinearly. It's displacement x at time t is given by x^2=at^2+b where a and b are constant. Its accelerations at time t is proportional to (a) 1/x^3

Hashim
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago

We are given that the displacement of the point is given by the following relation

 X2 = 1 + T2  (1)  

we know that the acceleration is the double derivative of the displacement. Hence it will be given by d2X / dT2

 from equation (1), we get

 X = (1 + T21/2  

differentiating with respect to T, we get

 dX / dT = 1/2 ( 1 + T2-1/2 x 2T

 = T ( 1 + T2-1/2

So 

 d2X / dT2 = 1/ ( 1 + T2)1/2 - T2 / (1+T23/2

But

 ( 1+ T2)1/2 = X  [ from (1)]

Thus 

 d2X / dT2 =  [1/ X ] - [(X2 -1) / X3]

 = [X2 -X2 +1] / X3

 = 1 / X3